OpenStudy (anonymous):

I am trying to find the total surface area of cones. I am having problems with finding the correct solution when the anwer is supposed to be 189 piecmtothe 2nd. The circumference is 18cm and the slant height is 12cm can someone show me how to work it out to get that answer?

OpenStudy (radar):

Insure that you are using an equation that is using "slant" heighth rather then perpendicular height, or calculate perpendicular height and us regular equation.

OpenStudy (anonymous):

Lateral Area of a right cone: ½ perimeter × slant height. I am not getting 189 \pi using this formula, I am getting 108 \pi

OpenStudy (anonymous):

I am trying to find the total surface area for the cylinder that has a diameter of 24m and height of 4m. the answer is 384msquared. How do I work this problem out.

OpenStudy (anonymous):

\[24 (\pi) 4 + 2 (12^2) \pi = 384 \pi \]

OpenStudy (anonymous):

The later area is the first part and the second part is top and bottom.

OpenStudy (anonymous):

lateral

OpenStudy (anonymous):

I figured it out. Thanks